# IRIS: An Immersive Robot Interaction System

### v2

@@ B-E1 Camera-to-Robot Base Transformation
B-E2 Point Cloud Processing Pipeline
@@ B-E3 XR Visualization and Teleoperation
B-E4 Calibration for XR-Robot Alignment
@@
B-E5 Future Extensions: Multi-Camera and ICP-Based Fusion
@@
B-F Affiliated Monitoring Tools
The dashboard depicted in 7 is designed to complement the IRIS Framework by providing streamlined access to the functionality offered by XR devices. Additionally, it facilitates the rendering of the scene to a web-based canvas, enabling real-time viewing of the scene as it is being streamed.
Due to the limitations of web services, which cannot directly interface with underlying protocols such as UDP and ZMQ, this website is unable to directly connect to the IRIS server. To bridge this gap, a Python-based endpoint is deployed, which hosts the website, relays the rendering data to the web application, and forwards user commands from the interface back to the server.
Communication between the endpoint and the website is facilitated through a WebSocket connection. This setup allows the system to support multiple instances of the dashboard simultaneously, ensuring scalability. The scene rendering itself is accomplished using Three.js, a powerful JavaScript library that renders to an HTML5 canvas object through WebGL. The Python endpoint is responsible for converting the IRIS scene format into commands, which are then transmitted over the WebSocket connection. These commands are subsequently translated into high-level Three.js function calls to render the scene.
This high-performance rendering system makes it possible to display even highly complex scenes on lower-end machines. Notably, only the server hosting the web application needs to install the necessary software, enabling all connected devices to access the interface without requiring any additional installations.
To minimize latency between the initial connection and the start of rendering, primitive objects are placed in the scene immediately upon receipt by the endpoint. Larger resources, such as meshes, materials, and textures, are streamed into the scene asynchronously to prevent delays or interruptions in the rendering process.
B-G Interaction Data
Here is an example of using ScenePub API to retrieve the data from Meta Quest 3 and HoloLen 2.
1# import device from scenepub
2from scenepub.xr_device.meta_quest3 import MetaQuest3
3from scenepub.xr_device.hololens2 import HoloLens2
4# do some thing about simulation
5mq3_player = MetaQuest3(mq3_device_name: str)
6hl2_player = HoloLens2(hl2_device_name: str)
7# running simulation
8while True:
9 # running simulation step logic
10 mc_data = mq3_player.get_mc_data()
11 hand_data = mq3_player.get_hand_data()
12 gaze_data = mq3_player.get_gaze_data()
13 # processing input data from Meta Quest3
14 hl2_hand_data = hl2_player.get_hand_data()
15 hl2_gaze_data = hl2_player.get_gaze_data()

### v3

@@ B.5.1 Camera-to-Robot Base Transformation
B.5.2 Point Cloud Processing Pipeline
@@ B.5.3 XR Visualization and Teleoperation
B.5.4 Calibration for XR-Robot Alignment
@@
B.5.5 Future Extensions: Multi-Camera and ICP-Based Fusion
@@
Appendix C Intuitive Robot Control Interface
C.1 Kinesthetic Teaching
Kinesthetic teaching is an intuitive interface that allows users to control robots by physically moving the real robot [71, 72, 7]. The real robot transmits joint positions and velocities in real time to the controlled robot, which can be either a virtual or another physical robot. In both real-world and simulation data collection, the key aspect of kinesthetic teaching is ensuring alignment between the real robot and its virtual counterpart in the XR headsets. By utilizing Spatial Anchors (C.4), the virtual robot in the XR headsets can be perfectly aligned with the real robot, which provide users with an intuitive and immersive experience.
C.2 Motion Controller
C.3 Affiliated Monitor Tools
The extensibility of IRIS opens up numerous possibilities for creating new applications. IRIS includes a web-based monitoring tool for managing all XR headsets. This tool allows users to easily start and stop alignment processes, as well as rename devices. Additionally, the tool supports real-time scene visualization using three.js [76], using the unified scene specification. An example screenshot is shown in Fig. 15.
C.4 Spatial Anchor
A spatial anchor serves as a reference point in a 3D environment to accurately position virtual objects within physical space. It enables virtual objects to maintain their position, orientation, and alignment, even as users move around or leave and return to the environment. IRIS utilizes either QR codes or motion controllers (currently only implemented for the Meta Quest 3) as spatial anchors. This approach allows for the alignment of augmented scenes across multiple headsets in the real world, making it appear as though all users are sharing the same scene. Additionally, this method can synchronize a virtual robot with its real-world counterpart in the Kinesthetic Teaching (C.1) interface and facilitate multiple-collector view alignment. Fig. 16 shows the alignment between real robot and virtual robot.
Figure 16: Alignment between real robot and virtual robot with a spatial anchor. It is used for controlling virtual robot by real world kinesthetic teaching or scene sharing among multiple users.
Appendix D Experiment Details
D.1 User Study
Refer to caption Refer to caption Refer to caption Refer to caption
(a1) (b1) (c1) (d1)
Refer to caption Refer to caption Refer to caption Refer to caption
(a2) (b2) (c2) (d2)
In this section, we provide more details of the user study we conducted, which assesses different data collection interfaces. As mentioned in the main paper, tasks from the LIBERO benchmark [45] were chosen due to their diversity in various movement patterns. Four representative tasks (see Fig. 17) were selected in the dimension of translation, rotation, and compound movement:
    Task 1: close the microwave
    Task 2: turn off the stove
    Task 3: pick up the book in the middle and place it on the cabinet shelf
    Task 4: turn on the stove and put the frying pan on it.
For the objective metrics, Table 2 presents the success rates for each interface across the tasks. A trial was considered unsuccessful if the participant failed to complete the task or exceeded the time limit. The time limits were set based on task difficulty: 20 seconds for Task 1 and Task 2, 30 seconds for Task 3, and 40 seconds for Task 4. The data shows a success rate of over $90\%$ across all four tasks when using the KT and MC interfaces from IRIS. In contrast, the Keyboard and 3D Mouse methods from LIBERO often resulted in failures. For example, the 3D Mouse interface achieved only a $37.5\%$ success rate on Task 3.
As is clear in Fig. 6, the results of this user study show that IRIS received higher scores than the baseline interfaces across both objective and subjective metrics. This indicates that the system offers a more intuitive and efficient approach for data collection.
D.2 Deformable Object Data Collection
Demonstrations involving deformable objects and cloth as seen in Fig. 18 can be recorded when using IsaacLab for physics simulation. Each point of a simulated deformable object is transmitted to IRIS. This allows for an accurate representation of object deformation during simulation. Transmitting each point individually adds a performance penalty compared to rigid objects. IRIS provides hand-tracking information that is used to control simulated robots and manipulate objects. The success criteria for the three example tasks as seen in Sec. 7 are as follows. Fold Cloth: The distance between two of the corner particles is below a threshold. This allows for folding either front-to-back or side-to-side. Lift Teddy: The center of the teddy must be lifted above a certain threshold. Stow Teddy: The center of the teddy must be inside the box and below a certain threshold. The initial object positions are chosen randomly without rotation. The box position for the Stow Teddy task is fixed. All example policies are using a U-Net diffusion model with reduced down dimension sizes of $[256,512,1024]$, and a diffusion step embedding dimension size of $256$. Other parameters are at their default values. The models are trained for 100 to 120 epochs. Observations are robot EEF information, and a mix of depth and image data. The last two observations are stacked and used as input. The models predict 16 actions at a time. Between 60 and 90 demonstrations are used for training.
Figure 18: Views wearing a VR headset and collecting task demonstrations using IRIS with IsaacLab. Tasks from left to right are: folding a cloth in half, lifting a deformable teddy, and stowing a deformable teddy in a slightly undersized box.
D.3 High Dynamic Data Collection
Traditional data collection methods struggle to acquire successful demonstrations in dynamic tasks that require high action timeliness and precision. To validate the practicality and effectiveness of IRIS, we applied it to a custom-designed table tennis task in MuJoCo simulation environment. IRIS was used to collect high-frequency strike data, and the collected demonstrations were evaluated using state-of-the-art imitation learning algorithms. In the task, a human demonstrator controls a virtual bat via a handheld controller to return randomly generated incoming balls, closely mimicking real bat manipulation. The system captures 73 around 5-second demonstrations at 100 Hz, where each demonstration movement consists of a forward stroke and natural backward release. The observation includes bat proprioceptive state and dual camera images, and the action is the desired bat position and orientation in task space. Fig. 8(a) shows the performance of imitation learning models trained on this data, using ball interception rate and successful return rate as evaluation metrics. The trained policy should first reach the incoming ball with natural stroke movement as human players and then hit the ball to opposite side table.
D.4 Real Robot Experiment
Figure 19: Two real-world manipulation tasks: Cup Inserting and Picking Up Lego.
To evaluate the real-world applicability of IRIS, we conducted a comparative study against a standard teleoperation baseline (Tele-Op) using two real-world manipulation tasks: Cup Inserting and Picking Up Lego (Fig. 19). For each task, 30 demonstrations were collected using both IRIS and Tele-Op, under identical experimental conditions. These datasets were then used to train two separate BC-Transformer policies, using consistent hyperparameters. We assessed both the data collection success rate (i.e., the percentage of successful demonstrations during collection) and the policy success rate (i.e., performance of trained policies in executing the tasks). As illustrated in Fig. 8(b), IRIS not only achieved a higher success rate during demonstration collection, but also resulted in better-performing policies compared to those trained on Tele-Op data. These results indicate that IRIS facilitates higher-quality real-world data collection, which in turn leads to more effective imitation learning outcomes.
Appendix E System Performance Analysis
The performance analysis focuses on two key aspects: network latency and headset FPS (frames per second). Since IRIS uses asynchronous bidirectional data transfer, network latency is low, averaging around 20-30 ms. Transmission bandwidth depends on Wi-Fi capacity. Even for large scenes from RoboCasa, which contain over 200 MB of compressed assets, it takes no more than 5 seconds to transfer and generate a full scene with more than 300 objects. For real robot teleoperation, the system handles point cloud data efficiently, achieving a transmission speed of 10,000 points at 60 Hz—exceeding the camera’s frame rate.